a populatpion of 1,000 students has an average weekly allowance of =350 and standard deviation of =56.13 .what is the probability that a random sample of size =30 will have an average weekly allowance between 335 and 360?


1
Expert's answer
2022-02-21T16:59:03-0500

Let "X=" an average weekly allowance: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=350, \sigma=56.13, n=30"


"P(335<X<360)=P(X<360)-P(X\le335)"

"=P(Z<\dfrac{360-350}{56.13/\sqrt{30}})-P(Z\le\dfrac{335-350}{56.13/\sqrt{30}})"

"\approx P(Z<0.975811)-P(Z\le-1.463716)"

"\approx 0.83542086-0.07163577"

"\approx 0.7638"